TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Create new blog entry for Point 8

Created by Main.AntonioVergara on 2007-07-02 - 22:12

Powering Tests Sector 7-8 - Report 02.07.2007

Microsoft Excel Spreadsheet PlanningW27_020707.xls:

Test Report from Monday 02.07.2007

* RQF.A78, RQF.A78 and RB.A78: Stability tests completed.

* RCBCH5.L8B1 stability test completed.

* 60A Closed Orbit Correctors: PNO completed for all the circuits released by ElQA. Tests OK except for two natural quenches in the circuits: RCBH28.L8B2 and RCBH25.L8B1

* 600A Correctors:

  • PCC and PIC2 completed for all the circuits available in DFBAO except RQTD.A78B1 and RQTD.A78B2
  • PCS completed successfully for RQTF.A78B1 and RQTF.A78B2
  • RCS.A78B2 - PCS.2 (50A) PCS.3 (100A) - Test OK
  • RCS.A78B1 - PCS.2 failed due to inversed polarity

Test Plan for Tuesday 03.07.2007

See file attached

Edit, attach file or add comment

Created by Main.BorisBellesia on 2007-07-02 - 12:11

Powering Test Sector 78 - 2nd July 2007

Microsoft Excel Spreadsheet PlanningW27_020707_.xls:


Edit, attach file or add comment

Created by Main.JacekSzkutnik on 2007-07-02 - 11:04

Cooldown Sector 78

We are now in the last week before the warm-up of sector 78 for the repair and connection of the inner triplet magnets and some consolidation work which is expected to last six weeks.

During the past week, the commissioning of sector 78 has continued on several fronts: the main magnet circuits, the 600 A corrector circuits, the 60A dipole correctors as well as the matching section magnets. This activity was made possible thanks to the very stable operation of the cryogenic system and of the utilities almost throughout the week. Unfortunately starting from Thursday morning two independent faults (return heater, 24V power supply) in the cryogenic system compromised the powering tests at 1.9K for the rest of the week.

After the two quadrupole circuits, which were both powered to 6.5 kA, the dipole circuit was successfully taken to 2 kA on Wednesday (June 27th) morning. The first cycle is shown below. This cycle ends with a discharge of the circuit in the dump resistor.

The in depth analysis of the data collected during these runs is ongoing to validate the correct operation of all the equipment.

The commissioning of the 600A circuit could finally start after the careful tuning of the quench detection system and the successful interlock test with the power converters connected to the magnets. The quench detection system is set to run with very low detection thresholds. The tests on these circuits are aimed at validating the intermagnet splices some of which could have been executed not to the required standards. These tests will continue this week.

Although not at full thrust, the commissioning of the 60 A orbit corrector circuits is progressing. In addition to all this, detailed studies on many circuits were performed to better identify the powering parameters of these circuits.

Studies to better understand the coupling between the Q4 and the D2 magnets did not succeed to reproduce the phenomenon which was initially observed despite all the efforts.

The preparation for the cool down of sector 45 is well advanced: the electrical quality assurance identified and localized a short in the dipole circuit which was very quickly repaired. A mechanical non conformity was also detected during the flushing: it concerned the swap of two lines in the interconnect between the low and high current module of the DFBAH. This was also quickly repaired and the conditioning of the lines continued over the weekend.

The puppets of the Joule-Thomson CV910 were also replaced and the cooldown is expected to start on Wednesday July 4th.

PNG Cooldown_78_Jul02_1.PNG:

PNG Cooldown_78_Jul02_2.PNG:


Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-07-02 - JacekSzkutnik
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login